# Statistics within UV–Visible Absorption Spectrum of Ethanolic Azobenzene

### v2

@@ Abstract
We report a statistical simulation of the UV–visible absorption spectrum of trans-azobenzene in ethanol solution. Due to intermolecular interactions, the used explicit solvent environment necessitates accounting for numerous transitions for a spectrum covering the two energetically lowest lines, S1 and S2. Furthermore, the spectrum manifests vast variation as a function of the underlying local structure, in conjunction with previous observations for spectra of liquids in the X-ray regime. We disentangle the complex structure–spectrum relationship using a machine learning-based method known as the emulator-based component analysis. This structural decomposition outperforms commonly used principal component analysis in explained spectral variation and reveals a small subset of latent structural variables responsible for the total spectral variance. Among other structural characteristics, blueshifting of the S2 peak occurs with fewer hydrogen bonds with the ethanol solvent, and a contracted N=N bond within the C–N=N–C bridge. The observed structural dependence of the absorption spectrum thus implies an overrepresentation of certain structural classes after a photoexcitation, potentially significant for the subsequent nuclear dynamics, photophysics, and photochemistry.

### v3

# Structural Decomposition of UV–Visible Spectral Variation: Azobenzene in Ethanol Solution

@@ Abstract
We present a structural interpretation of statistical variability in simulated liquid-phase UV–visible absorption spectra. We analyze the significant variation of the spectral response, caused by structural variation within the ensemble, using a response‑targeted method known as emulator‑based component analysis. In the high-dimensional input space, the method identifies a subspace of a few dimensions that accounts for most spectral variance. The resulting decomposition reveals the spectrally decisive structural features and filters out the irrelevant ones. For our test case, the ethanolic trans-azobenzene, the analysis implies an overrepresentation of certain structural characteristics following a photoexcitation at a given wavelength, potentially significant for the subsequent nuclear dynamics, photophysics, and photochemistry.